OpenStudy (anonymous):

y=3x-1 y=2x+8 solve for y and x

OpenStudy (anonymous):

First step, take y=3x-1 and plug it into the other one and then solve for x

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

sorry, im pretty dumb in math :$

OpenStudy (anonymous):

Plug it in, substitute the bottom equation becomes 3x-1=2x+8, then solve for x

OpenStudy (anonymous):

-2x+3x=8+1 x=9 is that right?

OpenStudy (anonymous):

and to find y all you do is just sub x back into any of the equations 3(9)-1 = y y=26

OpenStudy (anonymous):

Good job, you now know x. Now plug it in the next one and find y

OpenStudy (anonymous):

-2y+5y=8-2 3y=6 y=2 is that right?

OpenStudy (anonymous):

where did you get -2y+5y from?

OpenStudy (anonymous):

thats just another equation

OpenStudy (anonymous):

oh so different from the previous one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohhh okk yep that's correct then! =] so do you understand how to do it now?

OpenStudy (anonymous):

yeah ! thanks lol =)

OpenStudy (anonymous):

no worries